OpenStudy (anonymous):

A student uses the substitution method to solve the system of equations below algebraically. 2x - y = 5 3x + 2y = -3 Which equivalent equation could be used? 3x + 2(2x - 5) = -3 3x + 2(5 - 2x) = -3 3(y + 5/2) + 2y = -3 3(5/2 - y) + 2y = -3

OpenStudy (anonymous):

2x-y = 5. Now shuffle this equation to get the value of y: 2x+5 =y Substitute this (2x+5) for y in the 2nd equation, and you will have your answer.

OpenStudy (anonymous):

@livingmaybe Ok So the answer would be A?

OpenStudy (anthonyym):

No. You have 2x - y = 5 Add y. 2x = 5 + y Subtract 5 2x - 5 = y

OpenStudy (anonymous):

Sorry made a typo should have been y = 2x - 5. Yes A is right.

OpenStudy (anonymous):

Great Thank You @livingmaybe

OpenStudy (anthonyym):

A would be correct